The demand function of a product is given as P = 70 − 3Q, where P and Q are the price and quantity, respectively. Determine the price elasticity of demand if the price is equal to P = 10. [1] −1,167 [2] −0,960 [3] −0,800 [4] −0,167


1
Expert's answer
2021-05-25T08:38:16-0400

Price elasticity of demand:

"e_p=\frac{dQ}{dP}\cdot\frac{P}{Q}"

"Q=(70-P)\/3=(70-10)\/3=20"

"\frac{dQ}{dP}=-\frac{1}{3}"


"e_p=-\frac{1}{3}\cdot\frac{10}{20}=-0.167"

Answer: [4]